OpenStudy (anonymous):

Desperately need help with Langrangian optimization with inequality constraint!!! Please, please, please!! max (x,y,z) 12(x-y-z)-3x^2-4y^2-6z^2 s.t. x^2+y^2+z^2=<400 L 12(x-y-z)-3x^2-4y^2-6z^2-(lambda)x^2+y^2+z^2-400 dL/dx = 12-6x-(lambda)2x = 0 -> 12-6x=2x(lambda) dL/dy = 12-8y-(lambda)2y = 0 -> 12-8y=2y(lambda) dL/dz = 12-12z-(lambda)2z = 0 -> 12-12z=2z(lambda) dL/dlambda = x^2+y^2+z^2-400 = 0

OpenStudy (zarkon):

use standard methods to find local max/mins then reject any that don't satisfy \[x^2+y^2+z^2\le400\] then use Lagrange multipliers to check the boundary \[x^2+y^2+z^2=400\]

OpenStudy (anonymous):

I've gotten to the point of setting up the Lagrangian, but I can't figure out the solutions.

OpenStudy (zarkon):

why don't you show what you have so far

OpenStudy (anonymous):

Good point. I'll edit the question. Thank you!

OpenStudy (zarkon):

your dL/dy and dL/dz are a little off

OpenStudy (zarkon):

\[\frac{\partial }{\partial y}(12(x-y-z))=-12\]

OpenStudy (anonymous):

You take the partial derivative. X and Z drop off.

OpenStudy (zarkon):

\[\frac{\partial }{\partial y}(12(x-y-z))=\frac{\partial }{\partial y}(12x-12y-12z)\] \[=\frac{\partial }{\partial y}(12x)-\frac{\partial }{\partial y}(12y)-\frac{\partial }{\partial y}(12z)=0-12-0=-12\]